OpenStudy (anonymous):

what is the simplest form of sqrt[5]{x} times sqrt[5]{x} times sqrt[5]{x} times sqrt[5]{x}

OpenStudy (anonymous):

\[\sqrt[5]{x} \times \sqrt[5]{x} \times \sqrt[5]{x} \times \sqrt[5]{x} \]

OpenStudy (campbell_st):

well if you multiply the 1st 2 terms you get \[\sqrt[5]{x}\times \sqrt[5]{x} = \sqrt[5]{x^2}\] apply this rule to your problem... the 5th root and x don't change... its only the power of x that changes

OpenStudy (triciaal):